OpenStudy (precal):

Identify the increasing and decreasing intervals and identify the relative extrema.

OpenStudy (precal):

\[f(\theta)=\theta+2\sin\] on \[\left( 0,2 \pi \right)\]

OpenStudy (precal):

I took the first derivative and set it equal to zero to solve it

OpenStudy (precal):

\[f ' =1+2\cos \theta \]

OpenStudy (precal):

\[\theta=\frac{ 11 \pi }{ 6},\frac{ 7 \pi }{ 6}\]

OpenStudy (precal):

am I overlooking another solution because when I did a sign analysis, I notice that between these two values and using the unit circle cosine is negative and positive in that region

OpenStudy (precal):



zepdrix (zepdrix):

\(\Large\rm f(\theta)=\theta+2\sin\theta\) Your derivative looks good, \(\Large\rm f'(\theta)=1+2\cos\theta\) Your critical points do not. That's probably what is giving you trouble. Cosine is 1/2 at ummm pi/3. So our negative 1/2 will be at 2pi/3 and 4pi/3.

OpenStudy (precal):

yes thanks so much that is my issue